OpenStudy (anonymous):

put y= 4x^2 + 7x in vertex form. ( y= a(x-h)^2 + k I know that h is -7/8 but I can't find k.

jimthompson5910 (jim_thompson5910):

a = 4 since the leading coefficient is 4 y= a(x-h)^2 + k y= 4(x-(-7/8))^2 + k y= 4(x+7/8)^2 + k 0= 4(0+7/8)^2 + k ... plug in (0,0) since this point lies on y= 4x^2 + 7x 0 = 4(7/8)^2 + k 0 = 4(49/64) + k 0 = 49/16 + k keep going to solve for k

OpenStudy (anonymous):

So k is -49/16?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

yw